# Lightweight Transformer via Unrolling of Mixed Graph Algorithms for Traffic Forecast

### v1

@@ 6 Conclusion
To forecast city traffic, we unroll a mixed-graph optimization algorithm into a lightweight transformer-like neural net, where an undirected graph models spatial correlations and a directed graph models temporal relationships. We show that the two graph learning modules are akin to the classical self-attention mechanism. We design new $\ell_{2}$ and $\ell_{1}$-norm regularizers to quantify and promote signal smoothness on directed graphs. We interpret the derived processing operations from a constructed ADMM algorithm as low-pass filters. Experiments demonstrate competitive prediction performance at drastically reduced parameter counts. One limitation of our approach is that our computed Mahalanobis distances $\text{d}^{u}(i,j)$ and $\text{d}^{d}(i,j)$ for respective undirected and directed graphs are non-negative (due to learned PSD metric matrices ${\mathbf{M}}$ and ${\mathbf{P}}$), whereas affinity $e(i,j)$ in traditional self-attention can be negative. We will study extensions to signed distances as future work.
Appendix A Proof of Theorem 3.1
@@
As an illustration, consider a 4-node directed line graph ${\mathcal{G}}^{d}$ with edge weights $1$; see Fig. 2(a) for an illustration. We see that the symmetrized directed graph Laplacian ${\mathcal{L}}^{d}_{r}$ defaults to the graph Laplacian ${\mathbf{L}}^{u}$ for a 4-node undirected line graph ${\mathcal{G}}^{u}$ (Fig. 2(b)):
@@
Appendix B Interpreting Linear System (10)
@@
${\mathbf{r}}^{\tau}\triangleq\small({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\boldsymbol{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}.$ (54)
Solution ${\mathbf{x}}^{\tau+1}$ to linear system (10) can now be written as
@@
Define ~$\tilde{{\mathcal{L}}}_{r}^{d}\triangleq{\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}$. Note that ${\mathbf{H}}^{\top}{\mathbf{H}}$ is a diagonal matrix with zeros and ones (corresponding to chosen sample indices) along its diagonal, and thus ~$\tilde{{\mathcal{L}}}^{d}_{r}$ is a scaled variant of ${\mathcal{L}}^{d}_{r}$ with the addition of selected self-loops of weight $1$. Given that ${\mathcal{L}}^{d}_{r}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$ is real and symmetric, ~$\tilde{{\mathcal{L}}}^{d}_{r}$ is also real and symmetric. By eigen-decomposing ~$\tilde{{\mathcal{L}}}^{d}_{r}=\tilde{{\mathbf{U}}}\mathrm{diag}(\{\tilde{\xi}_{k}\})\tilde{{\mathbf{U}}}^{\top}$, we can rewrite ${\mathbf{x}}^{\tau+1}$ as
@@
Appendix C Deriving Solution to Optimization (15)
The optimization for ${\boldsymbol{\phi}}$ in objective (15) can be performed entry-by-entry. Specifically, for the $i$-th entry $\phi_{i}$:
@@ Appendix C Deriving Solution to Optimization (15)
Alternative solution with subgradients: Denote $g_{1}({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},g_{2}({\boldsymbol{\phi}})=({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|_{2}^{2}$, the subgradient of $g({\boldsymbol{\phi}})$ is
$\partial g({\boldsymbol{\phi}})=\{{\mathbf{g}}+\nabla g_{2}({\boldsymbol{\phi}})|{\mathbf{g}}\in\partial g_{1}({\boldsymbol{\phi}})\}=\{\mu_{d,1}\sigma+{\boldsymbol{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})|\sigma\in\partial(\|{\boldsymbol{\phi}}\|_{1})\},$ (61)
@@
Appendix D Parameterizing Conjugated Gradient Method in Solving Linear Equations
@@
3:${\mathbf{r}}_{0}\leftarrow{\mathbf{b}}-{\mathbf{A}}{\mathbf{x}}_{0}$triangleright Initial residual
4:${\mathbf{p}}_{0}\leftarrow{\mathbf{r}}_{0}$
5:$k\leftarrow 0$
@@
9:     ${\mathbf{r}}_{k+1}\leftarrow{\mathbf{r}}_{k}-\alpha_{k}{\mathbf{A}}{\mathbf{p}}_{k}$triangleright Dual residual
@@
Appendix E Model Setup
E.1 ADMM Blocks Setup
We initialize $\mu_{u},\mu_{d,1},\mu_{d,2}$ with 3, and the $\rho,\rho_{u},\rho_{d}$ with $\sqrt{N/(T+S+1)}$, where $N$ is the number of sensors and $T+S+1$ is the full length of the recovered signal (both the observed part and the part to be predicted). Those parameters are trained every ADMM layer. We initialize the CGD parameters (${\boldsymbol{\alpha}}$, ${\boldsymbol{\beta}}$)s as 0.08, and tune them for every CGD iteration. We clamp each element of the CGD step size ${\boldsymbol{\alpha}}$ to [0, 0.8] for stability. Each element of the momentum term coefficient ${\boldsymbol{\beta}}$ is clamped to be non-negative.
E.2 Graph Learning Modules Setup
@@
E.3 Feature Extractors Setup
@@
(b) A layer of eature extractor.
@@ E.3.1 Input Embeddings
where $\text{Var}(\text{d}^{S})$ denotes the variance of the physical distances. We use Laplacian embedding supposed by Dwivedi and Bresson [38], which uses the eigenvectors of the $k^{S}$-smallest non-zero eigenvalues of the Laplacian matrix ${\mathbf{L}}^{s}={\mathbf{D}}^{s}-{\mathbf{W}}^{s}$ to generate $k$-dimensional embeddings. Our experiments set $k^{S}=5$.
For temporal embeddings, we use the sine-cosine position embeddings[39]:
@@
We follow Feng and Tassiulas [9] to use real-time-stamp, time-in-day and day-in-week embeddings of dimensions 10, 6, 4.
@@
E.3.2 Feature Extraction from embeddings
We use a variant of GraphSAGE [40] to aggregate spatial features and introduce Temporal History Aggregation (THA) to aggregate spatial features. The GraphSAGE module aggregates the inputs of each node’s $k$-NNs with a simple and uniformed linear layer to generate $K$-dimension spatial features for each node of each time step. The THA module aggregates the embedded inputs of past $W$ time steps for each node with a uniformed linear layer, where $W$ is the size of the time window. For both GraphSAGE and THA modules, we pad zeros to inputs as placeholders, and use Swish [41] as acitvation functions with $\beta=0.8$.
E.4 Constructing Graphs ${\mathbf{W}}^{u,0},{\mathbf{W}}^{d,0}$ for the First ADMM Block
@@
We propose a simple module to make an initial guess of the signals: we modify the feature extractor described in Appendix E.3.2 to generate intermediate features from the embedded input of the observed part. Then we combine the feature and temporal dimensions, and use a simple linear layer with Swish activation to generate the initially extrapolated signal.
@@ E.5 Training Setup
Dataset PeMS03 PeMS04 METR-LA PEMS-BAY
$k$ 4 6 6 6
$W$ 6 6 6 4
Batch size 12 12 16 12
Appendix F Optimization Formulation and ADMM Algorithm in Ablation Studies
F.1 Optimization & Algorithm without DGTV Term
By removing the DGTV term $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ in (4), and introducing the auxiliary variables ${\mathbf{z}}_{u},{\mathbf{z}}_{d}$ as (9), we rewrite the unconstrained version of the optimization by augmented Lagrangian method as
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{z}}_{d}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$
$+({\boldsymbol{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|_{2}^{2}.$ (68)
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}+{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau},$ (69)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (70)
$\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{z}}_{d}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}.$ (71)
The updating of ${\boldsymbol{\gamma}}_{u}^{\tau},{\boldsymbol{\gamma}}_{d}^{\tau}$ follows (18).
F.2 Optimization & Algorithm without DGLR Term
By removing the DGLR term ${\mathbf{x}}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ in (4) and introduce only the auxiliary variables ${\boldsymbol{\phi}}$ in (5) and ${\mathbf{z}}_{u}$ in (9), we rewrite the unconstrained version of the optimization of ${\mathbf{x}}^{\tau+1}$ by augmented Lagrangian method as
$\min_{x,{\mathbf{z}}_{u}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{2}^{2}+({\boldsymbol{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$ (72)
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}$ $={\mathbf{H}}^{\top}{\mathbf{y}}+({\mathbf{L}}^{d}_{r})^{\top}(\frac{\gamma^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}$ (73)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}$ $=\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}$ (74)
The optimization of ${\boldsymbol{\phi}}^{\tau+1}$ is the same as (15), and the solution is the same as (16). The updating of ${\boldsymbol{\gamma}},{\boldsymbol{\gamma}}_{u}$ follows (17) and (18).
F.3 Algorithm Unrolling with an Undirected Temporal Graph
where $\mu_{u},\mu_{n}\in\mathbb{R}_{+}$ are weight parameters for the two GLR priors. The opitimization is similar to that in Appendix F.1. By a similar approach which introduces auxiliary variables ${\mathbf{z}}_{u}={\mathbf{x}},{\mathbf{z}}_{n}={\mathbf{x}}$, we optimize ${\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{n}$ as follows:
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\boldsymbol{\gamma}}_{u}^{\tau}+{\boldsymbol{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{n}}{2}{\mathbf{z}}_{n}^{\tau},$ (76)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (77)
$\left(\mu_{n}{\mathbf{L}}^{n}+\frac{\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{z}}_{n}^{\tau+1}=$ $\frac{{\boldsymbol{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{n}}{2}{\mathbf{x}}^{\tau+1}.$ (78)
The updating of ${\boldsymbol{\gamma}}_{u}^{\tau},{\boldsymbol{\gamma}}_{n}^{\tau}$ is also similar to (18) as follows:
${\boldsymbol{\gamma}}_{u}^{\tau+1}={\boldsymbol{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\quad{\boldsymbol{\gamma}}_{n}^{\tau+1}={\boldsymbol{\gamma}}_{n}^{\tau}+\rho_{n}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{n}^{\tau+1}).$ (79)
The undirected temporal graph ${\mathcal{G}}^{n}$ is learned by an UGL described in Section 4.3, and is initialized similar to the spatial UGL described in Appendix E.2. The initialization of $\mu_{n},\rho_{n}$ and the CGD parameters in solving (78) follows the setup in Appendix E.1.
Appendix G Further Ablation Studies
G.1 Sensitivity of $k$, $W$ in Graph Construction

### v2

@@ 6 Conclusion
To capture complex node-to-node relations in spatial-temporal data, we unroll a mixed-graph optimization algorithm into a lightweight transformer-like neural net, where an undirected graph models spatial correlations and a directed graph models temporal relationships. We show that the two graph learning modules play the role of self-attention, meaning that our unrolled network is a transformer. We design new $\ell_{2}$ and $\ell_{1}$-norm regularizers to quantify and promote signal smoothness on directed graphs. We interpret the derived processing operations from a designed ADMM algorithm as low-pass filters. Experiments demonstrate competitive prediction performance at drastically reduced parameter counts. One limitation of our approach is that our computed Mahalanobis distances $\text{d}^{u}(i,j)$ and $\text{d}^{d}(i,j)$ for respective undirected and directed graphs are non-negative (due to learned PSD metric matrices $\mathbf{M}$} and $\mathbf{P}$}), whereas affinity $e(i,j)$ in traditional self-attention can be negative. We will study extensions to signed distances and more complex graph modeling as future work.
Appendix A Related Works
A.1 Lightweight Learning Models
Given the importance of parameter reduction for increasingly large deep learning models, there exist various studies towards this common goal across application domains. For large language models (LLMs), smaller transformer models have been achieved via low-rank assumption (Hu et al., 2022) or parameter quantization (Frantar et al., 2023). For speech enhancement, Oostermeijer et al. (2021) proposes local causal self-attention, while Zhao & Madhu (2025) proposes parameter sharing across spectral and time dimensions. For low-light image enhancement, Brateanu et al. (2025) processes the Y-channel via convolution, pooling, and multi-head self-attention, separately from the U- and V-channels. For hyperspectral image classification, Sun et al. (2024) proposes a novel multiscale 3D-2D mixed CNN design. For cervical cancer detection, Pacal (2024) substitutes larger MBConv blocks in the MaxViT architecture with smaller ConvNeXtv2 blocks, and MLP blocks with GRN-based MLPs. Similar architecture miniaturization techniques have been proposed for edge devices (Mishra & Gupta, 2024) and resource-constrained environments (Liu et al., 2024).
For traffic prediction, there are practical use cases where parameter reduction is important. For example, drivers in a city predict local traffic conditions (in an area-specific neighborhood graph) on their memory-constrained mobile devices every five minutes to compute optimal routes to their destinations, based on updated congestion information shared in their local neighborhoods, instead of sending individual requests to the cloud and overwhelming the central server.
Thuc et al. (2024) achieves a lightweight model for image interpolation in a more principled manner: first design a linear-time optimization algorithm minimizing a chosen undirected graph smoothness prior (such as graph Laplacian regularizer (GLR) (Pang & Cheung, 2017) or graph total variation (GTV) (Bai et al., 2019)), then unroll its iterations into neural layers to compose a “white-box” transformer-like neural net for data-driven parameter learning. This algorithm unrolling approach results in a lightweight and mathematically interpretable network. Our current work differs from Thuc et al. (2024) in that we model the more complex node-to-node relationships in spatial-temporal data using a mixed graph: i) an undirected graph ${\mathcal{G}}^{u}$ to capture spatial correlations among geographically near stations, and ii) a directed graph ${\mathcal{G}}^{d}$ to capture sequential relationships along the temporal dimension. In particular, the notion of “frequencies” for directed graphs is not well understood in the graph signal processing (GSP) community. We are the first to define this notion and promote low-pass signal reconstruction via our designed $\ell_{2}$-norm directed graph Laplacian regularizer (DGLR) variational term in eq. 2 and $\ell_{1}$-norm directed graph total variation (DGTV) variational term in eq. 3, both derived from the graph shift operator (GSO) (Chen et al., 2015). To the best of our knowledge, we are the first define and promote smooth signals on a directed graph in a data-driven transformer setting.
A.2 Deep Models for Traffic Forecasting
Traffic forecasting problems contain spatial and temporal correlations between stations and time steps; thus, Graph Neural Network (GNN) is naturally employed in early research. Early research, such as STGCN (Yu et al., 2018), STSGCN (Song et al., 2020) and GGRNN (Ruiz et al., 2020), employs Graph Convolutional Networks (GCNs) to capture spatial relations and uses 1D convolutions and RNNs to process temporal relationships. ASTGNN (Ta et al., 2022), Graph WaveNet (Wu et al., 2019) and AGCRN (Bai et al., 2019) explored adaptive graph settings to capture potential non-local signal relations over the graph, trading in interpretability for improved performance. Graph Attention Networks (GATs) learn graph weights from embedded signals to enable flexible and data-targeted graph learning, typical examples include GMAN (Zheng et al., 2020) and ST-Wave (Fang et al., 2023). All GNNs above focus only on the spatial correlations, while the product graph (PG) is an intuitive model that describes local spatial-temporal relations at the same time. Einizade et al. (2024) presents a continuous product graph neural network for spatial-temporal forecasting, which achieves superior performance to the previous most-popular graph methods.
Transformers extend the powerful self-attention mechanism to all spatial-temporal data points, which enhances model performance with an extra-large number of parameters. Popular transformer-based models for traffic forecasting employ two transformers to process spatial and temporal dimensions separately (Xu et al., 2020; Chen et al., 2022; Huo et al., 2023), while state-of-the-art transformers introduce adaptive local attentions (Liu et al., 2023), time-delay (Jiang et al., 2023), or pre-training (Gao et al., 2024).
In recent years, simple linear models have been proposed to challenge the effectiveness and efficiency of heavy transformers. DLinear (Zeng et al., 2023) simply separates the input signals as trends and residuals, and employs 2 simple linear layers shared across nodes for prediction, but beats some of the transformer baselines in long-term forecasting. Other representative simple linear models include STID (Shao et al., 2022) and SimpleTM Chen et al. (2025).
As shown in Table 1, we select two most representative and well-performing models from each category as baselines for comparison.
Appendix B Proof of Theorem 3.1
@@ A.2 Deep Models for Traffic Forecasting
As an illustration, consider a 4-node directed line graph ${\mathcal{G}}^{d}$ with edge weights 11; see Fig. 2(a) for an illustration. We see that the symmetrized directed graph Laplacian ${\mathcal{L}}^{d}_{r}$ defaults to the graph Laplacian ${\mathbf{L}}^{u}$ for a 4-node undirected line graph ${\mathcal{G}}^{u}$ (Fig. 2(b)):
@@
Appendix C Interpreting Linear System eq. 10
@@
${\mathbf{r}}^{\tau}\triangleq({\mathbf{L}}^{d}_{r})^{\top}(\frac{{\bm{\gamma}}^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}-\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau}.$ (54)
Solution ${\mathbf{x}}^{\tau+1}$ to linear system eq. 10 can now be written as
@@
Define $\tilde{{\mathcal{L}}}_{r}^{d}\triangleq{\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}$. Note that ${\mathbf{H}}^{\top}{\mathbf{H}}$ is a diagonal matrix with zeros and ones (corresponding to chosen sample indices) along its diagonal, and thus $\tilde{{\mathcal{L}}}^{d}_{r}$ is a scaled variant of ${\mathcal{L}}^{d}_{r}$ with the addition of selected self-loops of weight $1$. Given that ${\mathcal{L}}^{d}_{r}=({\mathbf{L}}^{d}_{r})^{\top}{\mathbf{L}}^{d}_{r}$ is real and symmetric, $\tilde{{\mathcal{L}}}^{d}_{r}$ is also real and symmetric. By eigen-decomposing $\tilde{{\mathcal{L}}}^{d}_{r}=\tilde{{\mathbf{U}}}\mathrm{diag}(\{\tilde{\xi}_{k}\})\tilde{{\mathbf{U}}}^{\top}$, we can rewrite ${\mathbf{x}}^{\tau+1}$ as
@@
Appendix D Deriving Solution to Optimization eq. 15
The optimization for $\boldsymbol{\phi}$} in objective eq. 15 can be performed entry-by-entry. Specifically, for the $i$-th entry $\phi_{i}$:
@@
Alternative solution with subgradients: Denote $g_{1}({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1},g_{2}({\boldsymbol{\phi}})=({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|_{2}^{2}$, the subgradient of $g({\boldsymbol{\phi}})$ is
$\partial g({\boldsymbol{\phi}})=\{{\mathbf{g}}+\nabla g_{2}({\boldsymbol{\phi}})|{\mathbf{g}}\in\partial g_{1}({\boldsymbol{\phi}})\}=\{\mu_{d,1}\sigma+{\bm{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})|\sigma\in\partial(\|{\boldsymbol{\phi}}\|_{1})\},$ (61)
@@
Appendix E Parameterizing Conjugated Gradient Method in Solving Linear Equations
@@
[…]
@@
9:  ${\mathbf{r}}_{k+1}\leftarrow{\mathbf{r}}_{k}-\alpha_{k}{\mathbf{A}}{\mathbf{p}}_{k}$ $\triangleright$ Dual residual
@@
Appendix F Model Setup
F.1 ADMM Blocks Setup
We initialize $\mu_{u},\mu_{d,1},\mu_{d,2}$ with 3, and the $\rho,\rho_{u},\rho_{d}$ with $\sqrt{N/(T+S+1)}$, where $N$ is the number of sensors and $T+S+1$ is the full length of the recovered signal (both the observed part and the part to be predicted). Those parameters are trained every ADMM layer. We initialize the CGD parameters ($\bm{\alpha}$}, $\bm{\beta}$})s as 0.08, and tune them for every CGD iteration. We clamp each element of the CGD step size $\bm{\alpha}$} to [0, 0.8] for stability. Each element of the momentum term coefficient $\bm{\beta}$} is clamped to be non-negative.
F.2 Graph Learning Modules Setup
@@
F.3 Feature Extractors Setup
@@
(b) A layer of feature extractor.
@@ F.3.1 Input Embeddings
where $\text{Var}(\text{d}^{S})$ denotes the variance of the physical distances. We use Laplacian embedding supposed by Dwivedi & Bresson (2020), which uses the eigenvectors of the $k^{S}$-smallest non-zero eigenvalues of the Laplacian matrix ${\mathbf{L}}^{s}={\mathbf{D}}^{s}-{\mathbf{W}}^{s}$ to generate $k$-dimensional embeddings. Our experiments set $k^{S}=5$.
For temporal embeddings, we use the sine-cosine position embeddings (Vaswani et al., 2018):
@@
We follow Feng & Tassiulas (2022) to use real-time-stamp, time-in-day and day-in-week embeddings of dimensions 10, 6, 4.
@@
F.3.2 Feature Extraction from embeddings
We use a variant of GraphSAGE (Hamilton et al., 2017) to aggregate spatial features and introduce Temporal History Aggregation (THA) to aggregate spatial features. The GraphSAGE module aggregates the inputs of each node’s $k$-NNs with a simple and uniformed linear layer to generate $K$-dimension spatial features for each node of each time step. The THA module aggregates the embedded inputs of past $W$ time steps for each node with a uniformed linear layer, where $W$ is the size of the time window. For both GraphSAGE and THA modules, we pad zeros to inputs as placeholders, and use Swish (Ramachandran et al., 2017) as acitvation functions with $\beta=0.8$.
F.4 Constructing Graphs ${\mathbf{W}}^{u,0},{\mathbf{W}}^{d,0}$ for the First ADMM Block
@@
We propose a simple module to make an initial guess of the signals: we modify the feature extractor described in Appendix F.3.2 to generate intermediate features from the embedded input of the observed part. Then we combine the feature and temporal dimensions, and use a simple linear layer with Swish activation to generate the initially extrapolated signal.
@@ F.5 Training Setup
Dataset PeMS03 PeMS04 METR-LA
$k$ 4 6 6
$W$ 6 6 6
Batch size 12 12 16
Appendix G Training Performance w.r.t. Training Data Cost
Due to the parameter efficiency of our model, our unrolling model is also data-efficient under the same observation length. We trained our models and baselines with reduced training datasets by selecting sample windows from the sequence with larger strides of 5 and 10 in the PEMS03 dataset for 60-minute forecasting. The results below show that baseline models with large parameter counts suffer from steep performance degradation when the training data becomes scarce. In contrast, our model maintains stable performance under limited data conditions, demonstrating both lower data requirements and better generalizability. This makes it particularly suitable for deployment in data-scarce or resource-constrained environments.
Table 4: Performance under different sampling strides (metrics: MAE / RMSE / MAPE(%)) on 60-minute forecasting for PEMS03 dataset.
Model Stride = 3 (original) Stride = 5 Stride = 10
Ours 27.67 / 17.46 / 17.72 30.42 / 18.60 / 20.05 30.65 / 19.02 / 19.67
VAR 30.54 / 18.31 / 19.61 30.68 / 18.37 / 19.74 31.25 / 18.71 / 21.22
STSGCN 30.62 / 19.35 / 19.15 32.37 / 20.67 / 20.66 33.22 / 21.30 / 21.58
PDFormer 27.16 / 17.26 / 21.21 25.70 / 15.37 / 15.99 77.29 / 44.36 / 126.63
AGCRN 29.90 / 16.76 / 15.32 39.25 / 20.60 / 16.95 62.58 / 109.27 / 30.72
Appendix H Ablation Studies
H.1 Ablation Studies for $\ell_{2}$- and $\ell_{1}$- norms in Direct Graph Modeling
First, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, as well as the benefit of directed-graph modeling of temporal relations. We separately test the effects of DGLR and DGTV by removing the prior and unrolling the ADMM algorithm. We denote these models as “w/o DGTV” and “w/o DGLR”. To test the benefit of the directed-graph modeling, we change the temporal graph ${\mathcal{G}}^{d}$ into a fully undirected graph ${\mathcal{G}}^{n}$ while maintaining the connections, and replace the directed graph priors with a GLR prior ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ with the same coefficient $\mu_{n}=\mu_{d,2}$. We denote the undirected-time unrolling model as “UT”. Detailed derivations of the ADMM algorithm for all the above ablation studies are in Appendix I.
Dataset Mixed-Graph UT w/o DGTV w/o DGLR
METR-LA 12.34 / 5.18 / 11.80 12.51 / 5.42 / 12.11 12.36 / 5.40 / 12.31 12.41 / 5.35 / 12.20
H.2 Sensitivity of $k$, $W$ in Graph Construction